The sum of three times a number and 7 more than the number is the same as the difference between -11 and twice the number. What is the number?
I have tried:
3x+7=-11-2x
3x+2x=-11-7
x=-18
This is wrong according to my instructor. The answer is actually supposed to be -3. How?

Found 2 solutions by Earlsdon, Alan3354:
Answer by Earlsdon(6294)

Try this!
3x+(x+7) "The sum (+) of three times a number (3x) and seven more than the number (x+7)" is the same as:
-11-2x "the difference (-) between -11 and twice the number (2x)
Putting it al together, we get:
3x+x+7 = -11-2x Simplify.
4x+7 = -11-2x Add 2x to both sides.
6x+7 = -11 Subtract 7 from both sides.
6x = -18 Finally, divide both sides by 6.
x = -3 Voila!

Answer by Alan3354(69443)

The sum of three times a number and 7 more than the number is the same as the difference between -11 and twice the number. What is the number?
I have tried:
3x+7=-11-2x
3x+2x=-11-7
x=-18
This is wrong according to my instructor. The answer is actually supposed to be -3. How?
----------------
3 times the number is 3x.
7 more than the number is x+7.
----------------------------
3x + x+7 = -11-2x
4x+7 = -11-2x
6x = -18
x = -3
----------
Just some tricky wording. I don't see what that's supposed to accomplish.